The following table shows the distribution of the number of donations made by a class of students last month:
Number of donations made 1 2 3 4 5 6
Number of students 20 5 3 2 4 2
If a students is elected at random, find the probabilities of the following:
The student made more than 3 donations last month

Answer by math_tutor2020(3816)


Answer: 2/9

----------------------------------
Work Shown:

The given table is
x123456
y2053242

x = number of donations made
y = number of students

We have 2+4+2 = 8 students who made more than 3 donations. We find this by adding the values of y that correspond to x = 4, x = 5, x = 6; all of these x values are greater than 3 (to indicate more than 3 donations were made).

There are 20+5+3+2+4+2 = 36 students total.

The probability of getting a student who made more than three donations is therefore:
8/36 = (4*2)/(4*9) = 2/9